OpenStudy (anonymous):

Find the half-life of an element, which decays at a rate of 3.411% per day?

OpenStudy (anonymous):

\[1-.03411=0.96589\] solve \[(0.96589)^x=.5\] via the change of base formula \[x=\frac{\ln(.5)}{\ln(0.96589)}\] and a calculator

OpenStudy (anonymous):

Can you explain how you know to do that?

OpenStudy (anonymous):

oh how i know it? \[b^x=A\iff x=\frac{\ln(A)}{\ln(b)}\] it is called the "change of base formula" allows to you solve when the variable is in the exponent

OpenStudy (anonymous):

yeah i get that part. i just don't know why you took the decay rate and subtracted it from 1 and all that.

OpenStudy (anonymous):

and to decrease a number by \(3.411\%\) you multiply it by \[100\%-3.411\%=96.589\%=.96589\]

OpenStudy (anonymous):

and where did this formula come from: (0.96589)^x=.5?

OpenStudy (anonymous):

that is unless you were taught to use \[e^{-.03411t}\] but that is not really what it says

OpenStudy (anonymous):

half life means you end up with half of what you started with that is why i set it equal to \(\frac{1}{2}\) or \(.5\)

OpenStudy (anonymous):

all we were given for decay/growth was:\[q=q _{0}e ^{r*t}\]

OpenStudy (anonymous):

ok then go with \[\large e^{-.03411t}=.5\] so \[-.03411t=\ln(.5)\] or \[t=\frac{\ln(.5)}{-.03411}\]

OpenStudy (anonymous):

half life, set it equal to one half then solve for \(t\) as above

OpenStudy (anonymous):

that makes more sense. I tried solving for t but got a negative number which didn't look right. but i didn't set .03411 to be negative